An empty bottle of mass 3 kg is released from a submarine and rises with an

acceleration of 0.7 ms?. If the water causes a resistance of 0.5 N, the size of the

buoyancy force causing the bottle to rise is
Answer #1
3*9.8=29.4N 29.4+0.5=29.9N F-29.9=3x0.7 F=3x0.7+29.9 F=32N
answered by: anonymous